A Bibliography of Publications in Notices of the American Mathematical Society (1970–1979)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

15 August 2024 Version 1.00

Title word cross-reference

2 × 2 [Tau75]. p [EM75]. Q_{p} [EM75]. X_{m} + 1 [EM75]. -Adic [EM75]. 1973 [Tau74]. 20 [Tau74]. 25 [Tau78]. Adic [EM75]. Algebras [BU75]. AMS [Tau78, Tau74]. Answer [Tau78]. based [FK71]. Baumert [Tau78]. Calculus [BU75]. classes [Tau71].

Clenshaw [Rie71]. compounding [Leh70]. concerning [Tau71, Tau73]. Curtis [Rie71]. D [Tau78]. diagonal [Tau73]. Error [Rie71]. estimates [Rie71]. Expository [SG83]. Field [EM75]. fields [Tau75]. Halmos [SG83]. L [Tau78]. Mathematics [Hal74]. matrices [Tau71, Tau75]. matrix [Tau73]. means [Leh70]. minimization [FK71]. noncommuting [Tau75]. Norms [Tau75]. Notices [Tau78, Tau74].

October [Tau74]. off [Tau73].

P. [SG83]. part [Tau73]. Preliminary [BU75, Rie71]. Projective [BU75].

quadratic [Tau75]. quadrature [Rie71]. Query [Tau74, Tau78].

R [SG83]. rational [Tau75]. Recurrence [FK71]. relation [Tau75]. Relations [BU75, FK71]. remarks [Tau73]. Report [BU75, Rie71]. Response [Tau74]. Result [Tau71]. Roots [EM75].

Selecta [SG83]. Small [Tau74]. Some [Tau73].

Talk [Hal74]. their [Tau75].

Writing [SG83].

References

Bednarek:1975:PAC

[BU75] A. R. Bednarek and S. M. Ulam. Projective algebras and calculus of relations — preliminary report. Notices of the American Mathematical Society, 22(7):A729, November 1975. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Everett:1975:RXP

[EM75] C. J. Everett and N. Metropolis. On the roots of X_{m} + 1 in the p-adic field Q_{p}. Notices of the American Mathematical Society, 22(6):A619, October 1975. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Fredman:1971:RRB

[FK71] Michael L. Fredman and Donald E. Knuth. Recurrence relations based on minimization. Notices of the American Mathematical Society, 18(6):960, October 1971. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Halmos:1974:HTM

[Hal74] Paul R. Halmos. How to talk mathematics. Notices of the American Mathematical Society, 21(3):155–158, April 1974. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Reprinted in [SG83].

Lehmer:1970:CM

[Leh70] D. H. Lehmer. On compounding of means. Notices of the American Mathematical Society, 17(4):634–635, June 1970. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Riess:1971:EECa

[Rie71] R. D. Riess. Error estimates for Clenshaw–Curtis quadrature: preliminary report. Notices of the American Mathematical Society, 18(5):790– 791, August 1971. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https:/

/www.ams.org/journals/notices/
197108/197108FullIssue.pdf.

Sarason:1983:PRH

[SG83] Donald E. Sarason and Leonard Gillman, editors. P. R. Halmos. Selecta: Expository Writing. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. ISBN 0-387-90756-4. ???? pp.

Taussky:1971:RCC

[Tau71] Olga Taussky. Result concerning classes of matrices. Notices of the American Mathematical Society, 18 (6):943, October 1971. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Taussky:1973:SRC

[Tau73] Olga Taussky. Some remarks concerning off diagonal part of a matrix. Notices of the American Mathematical Society, 20(6):A567, October 1973. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Taussky:1974:RQS

[Tau74] Olga Taussky. Response to query 26 (Small, Notices AMS 20:307, October 1973). Notices of the American Mathematical Society, 21(3):159, April 1974. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Taussky:1975:NQF

[Tau75] Olga Taussky. Norms in quadratic fields and their relation to noncommuting 2 × 2 rational matrices. Notices of the American Mathematical Society, 22(2):A304, February 1975. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Taussky:1978:AQD

[Tau78] Olga Taussky. Answer to query 153 (L. D. Baumert, Notices AMS, 25:252 (1978)). Notices of the American Mathematical Society, 25(6):424, October 1978. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).